# Two-Edge Connectivity via Pac-Man Gluing

### v1

@@ 4 Canonical 2-Edge Cover
It was shown in [27] that a minimum triangle-free 2-edge cover can be computed in polynomial time.
Lemma 4 (Proposition 7 in [27]).
@@
Notice that for our input structured graph $G$, a 2-ECSS of $G$ is also a triangle-free 2-edge cover. Thus, if $H$ is a minimum triangle-free 2-edge cover of our input $G$, then $\lvert H\rvert\leq\mathsf{opt}(G)$. Roughly speaking, we will compute $H$ using the above lemma, and then transform $H$ into a 2-ECSS of $G$ by adding at most $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$ edges to it. Thus, we would get a 2-ECSS with at most $\nicefrac{{5}}{{4}}\cdot\mathsf{opt}(G)$, proving our main result. In reality, we might add some more edges and delete some edges, so that the net total of additional edges will be at most $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$.
We will first convert $H$ into a canonical triangle-free 2-edge cover without changing the number of edges in it. In order to define canonical we first need the following definitions pertaining to a 2-edge cover.
@@
Let $H$ be a $2$-edge cover of some graph. $H$ is bridgeless if it contains no bridges, i.e., all the components of $H$ are $2$-edge connected. A component of $H$ is complex if it contains a bridge. Maximal222w.r.t. the subgraph relationship $2$-edge-connected subgraphs of $H$ are blocks. A block $B$ of some complex component $C$ of $H$ is called pendant if $C\setminus V(B)$ is connected. Otherwise it is non-pendant. Note that by contracting each block of a complex component to a single vertex, we obtain a tree whose leaves correspond to pendant blocks.
@@ A $2$-edge cover $H$ is canonical if the following properties are satisfied:
  • For every complex component, each of its pendant blocks contains at least $6$ edges and each of its non-pendant blocks contains at least $4$ edges.
We establish the following lemma in Appendix B that shows how to convert a minimum triangle-free 2-edge cover of a structured graph into a minimum canonical 2-edge cover in polynomial time.
@@
lemmalemmaCanonicalMain Given a structured graph $G$, in polynomial time we can compute a canonical $2$-edge cover $H$ of $G$, where the number of edges in $H$ is at most (and hence equal to) the number of edges in a minimum triangle-free $2$-edge cover. This definition of canonical and the proof of the above lemma are similar to the ones in [18], except that now components with $7$ edges have to be a $\mathcal{C}_{7}$.
@@ 5 Bridge Covering and Credits
We now have a structured graph $G$ and a minimum canonical 2-edge cover $H$ of $G$ such that $|H|\leq\mathsf{opt}(G)$. We will show how to convert $H$ into a 2-ECSS of $G$ by adding a net total of at most $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$ edges to it. There are two reasons why $H$ is not already a 2-ECSS of $G$: First, $H$ might have multiple components. Second, some of these components might have bridges. We will first modify $H$ so that the resulting components do not have any bridges. Then, we will glue the various resulting 2EC components together to get a desired 2-ECSS, which we defer to the next section. Throughout this transformation, we will keep the property that the intermediate solution is a canonical 2-edge cover.
To keep track of how many edges we are adding, we will define a credit invariant that is satisfied throughout the transformation and a cost function. We have a total budget of $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$, which we will distribute all over $H$ as credits of different kinds. The cost of $H$ is defined keeping in mind that the cost of a single edge is 1. $\mathrm{cost}(H)$ is the total number of edges in $H$ plus the total credits in $H$. Thus, initially the $\mathrm{cost}(H)$ will be at most $\nicefrac{{5}}{{4}}\cdot\mathsf{opt}$. At each step of our transformation, we aim to ensure that the cost never increases, so that the final solution we obtain will have cost at most $\nicefrac{{5}}{{4}}\cdot\mathsf{opt}$, and we will be done. We now define our credit invariant and the cost function, which is also used in the next section.
@@
We first show that we can distribute our budget of $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$ over our minimum canonical triangle-free 2-edge cover to satisfy the credit invariant. In other words, we assume the credits are there as required by the credit invariant and show that $\mathrm{cost}(H)\leq\nicefrac{{5}}{{4}}\cdot\lvert H\rvert$.
Lemma 5.
Let $H$ be a canonical $2$-edge cover of a graph $G$. Then, we have $\mathrm{cost}(H)\leq\frac{5}{4}|H|$.
@@
We are now ready to state our main lemma for bridge covering, which essentially is already proved in [18]. We add some comments in Appendix C on how to adapt the proof to our setting.
@@
We will feed our minimum canonical 2-edge cover $H$ to the algorithm in the above lemma to obtain a bridgeless canonical 2-edge cover $H^{\prime}$ and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)\leq\nicefrac{{5}}{{4}}\cdot\mathsf{opt}(G)$. Each component of $H^{\prime}$ now is 2EC and either has $8$ or more edges, in which case it has a credit of $2$, otherwise it is a cycle with $i$ edges ($\mathcal{C}_{i}$) for $4\leq i\leq 7$ with $\nicefrac{{1}}{{4}}\cdot i$ credits. We will use these credits to glue the 2EC components of $H^{\prime}$ together to obtain a desired 2-ECSS of $G$, as explained in the next section.
@@ 6 Gluing
Using the above lemma, the proof of Theorem 1 follows easily: For a structured graph $G$, we output the 2-ECSS $H^{\prime}$ of $G$ obtained by applying the lemma above. We know $\lvert H^{\prime}\rvert+\mathrm{cr}(H^{\prime})=\mathrm{cost}(H^{\prime})\leq\frac{5}{4}\cdot\mathsf{opt}(G)+3$. Now $\mathrm{cr}(H^{\prime})=2$, as it will be a single 2EC component with more than $8$ edges. Thus, $\lvert H^{\prime}\rvert\leq\frac{5}{4}\cdot\mathsf{opt}(G)+1=(\frac{5}{4}+\frac{\varepsilon}{4})\cdot\mathsf{opt}(G)$, where the last inequality follows from $\mathsf{opt}(G)\geq\lvert V(G)\rvert\geq\frac{4}{\varepsilon}$, which holds for structured graphs $G$. Now, using the approximation preserving reduction, Definition 5, for general graphs $G$, we incur an additional loss of $\varepsilon$ to get an approximation ratio of $\frac{5}{4}+\frac{5}{4}\varepsilon=\frac{5}{4}+\varepsilon^{\prime}$, where $\varepsilon^{\prime}=\frac{5}{4}\varepsilon$, proving Theorem 1.
It remains to prove Lemma 6. To this end, we first introduce some definitions. In what follows, unless otherwise mentioned, $G$ is a structured graph and $H$ is a bridgeless canonical 2-edge cover of $G$.
@@
The component graph ^$\bm{\hat{}}{G}_{H}$ w.r.t $G$ and $H$ is the multigraph obtained from $G$ by contracting the vertices of each 2EC component of $H$ into a single node. We call the vertices of ^$\bm{\hat{}}{G}_{H}$ nodes to distinguish them from the vertices in $G$. There is a one-to-one correspondence between the edges of ^$\bm{\hat{}}{G}_{H}$ and $G$. Furthermore, there is a one-to-one correspondence between the nodes of ^$\bm{\hat{}}{G}_{H}$ and the components of $H$. We will denote the component corresponding to a node $A$ by $C_{A}$.
@@
A non-trivial segment $S$ of ^$\bm{\hat{}}{G}_{H}$ is a maximal 2-node-connected subgraph of ^$\bm{\hat{}}{G}_{H}$ consisting of at least 3 nodes. A trivial segment is a single node of ^$\bm{\hat{}}{G}_{H}$ that is not part of any non-trivial segment. For a segment $S$, $V(S)$ refers to the set of all the vertices in the various components corresponding to the nodes contained in $S$.
Notice that we can break the component graph into segments such that two segments can potentially overlap only at a cut node. This is easily achievable by identifying all the cut nodes of the component graph.
Furthermore, since a node $A$, that is a trivial segment, is not part of any cycles of length at least 3 in ^$\bm{\hat{}}{G}_{H}$ (otherwise $A$ would be part of a non-trivial segment), $A$ has to be either a cut node of ^$\bm{\hat{}}{G}_{H}$ or a pendant node (neighbor of exactly one other node) of ^$\bm{\hat{}}{G}_{H}$. In either case, we can apply the 3-matching lemma, Lemma 2, to argue that $V(C_{A})$ has a matching of size 3 going out to $V(C_{B})$, where node $B$ is a neighbor of node $A$ in ^$\bm{\hat{}}{G}_{H}$. The condition of the lemma is satisfied by partitioning $V(G)$ into two parts such that one part has vertices from $V(C_{A})$ and the other part has vertices from $V(C_{B})$ and the only edges going across the parts are between $V(C_{A})$ and $V(C_{B})$. We have the following propositions.
@@
Let $A$ be a trivial segment of ^$\bm{\hat{}}{G}_{H}$. Then, $A$ is either a cut node of ^$\bm{\hat{}}{G}_{H}$ or it is a pendant node of ^$\bm{\hat{}}{G}_{H}$.
@@
Let $A$ and $B$ be adjacent nodes in ^$\bm{\hat{}}{G}_{H}$ such that $A$ is a trivial segment. Then, there exists a matching of size $3$ going between $V(C_{A})$ and $V(C_{B})$ in $G$.
Having set up the requisite definitions, we now outline the steps involved in the gluing process. First, if $H$ does not contain a huge component, then we first glue a few components together, so that we obtain a canonical bridgeless 2-edge cover $H^{\prime}$ that has a huge component, while maintaining the credit invariant. We will show that such an $H^{\prime}$ can be obtained with $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)+3$. After this step, we assume $H$ has a huge component, an invariant that will be maintained at each iteration of our algorithm. Next, let $L$ be a huge component in $H$. We will distinguish two cases: either $L$ is a trivial segment of ^$\bm{\hat{}}{G}_{H}$, or it is part of a non-trivial segment of ^$\bm{\hat{}}{G}_{H}$. In either case we show that we can make progress, i.e., we can find in polynomial time a canonical bridgeless 2-edge cover $H^{\prime}$ that has fewer components than $H$ such that $H^{\prime}$ contains a huge component and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$. Exhaustively applying the above step will result in a single bridgeless 2EC component, a 2-ECSS of $G$. Thus, if we can establish the existence of these steps, Lemma 6 follows immediately. These steps are encapsulated in the following three lemmas.
@@
Let $C_{L}$ be a huge component such that $L$ is a trivial segment of ^$\bm{\hat{}}{G}_{H}$ and $H\neq C_{L}$. Then, in polynomial time we can compute a canonical bridgeless 2-edge cover $H^{\prime}$ of $G$ such that $H^{\prime}$ has fewer components than $H$, $H^{\prime}$ contains a huge component, and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$.
@@
Let $C_{L}$ be a huge component such that $L$ is part of a non-trivial segment of ^$\bm{\hat{}}{G}_{H}$. Then, in polynomial time we can compute a canonical bridgeless 2-edge cover $H^{\prime}$ of $H$ such that $H^{\prime}$ has fewer components than $H$, $H^{\prime}$ contains a huge component, and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$.
As explained above, the above three lemmas immediately imply Lemma 6. All that remains is to prove the above three lemmas, which we do in the next three subsections.
6.1 Proof of Lemma 9

### v2

@@ 4 Canonical 2-Edge Cover
It was shown in [KN23] that a minimum triangle-free 2-edge cover can be computed in polynomial time using a polynomial-time algorithm for computing a maximum triangle-free 2-matching, e.g. [Har24, Pal23].
Lemma 4 (Proposition 7 in [KN23]).
@@
Notice that for a structured graph $G$, a 2-ECSS of $G$ is also a triangle-free 2-edge cover because $G$ contains at least $\nicefrac{{4}}{{\varepsilon}}$ vertices, and thus, a 2-ECSS of $G$ cannot be a triangle. Thus, if $H$ is a minimum triangle-free 2-edge cover of our input $G$, then $\lvert H\rvert\leq\mathsf{opt}(G)$. Roughly speaking, we will compute $H$ using the above lemma, and then transform $H$ into a 2-ECSS of $G$ by adding at most $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$ edges to it. Thus, we would get a 2-ECSS with at most $\nicefrac{{5}}{{4}}\cdot\mathsf{opt}(G)$, proving our main result. In reality, we might add some more edges and delete some edges, so that the net total of additional edges will be at most $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$.
We will first convert $H$ into a canonical triangle-free 2-edge cover without changing the number of edges in it. To define canonical, we first need the following definitions pertaining to a 2-edge cover.
@@
Let $H$ be a $2$-edge cover of a structured graph $G$. We call $H$ bridgeless if it contains no bridges, i.e., all the components of $H$ are 2EC. A component of $H$ is complex if it contains a bridge. Given a complex component $C$, a maximal111w.r.t. the subgraph relationship 2EC subgraph of $H$ containing at least 3 (as $G$ has no parallel edges) vertices are blocks. Any vertex of $C$ not contained in a block is called lonely. A block $B$ of some complex component $C$ of $H$ is called pendant if $C\setminus V(B)$ is connected. Otherwise, it is non-pendant.
Note that by contracting each block of a complex component to a distinct node, we obtain a tree whose leaves correspond to pendant blocks. We are now ready to define canonical 2-edge covers.
@@ A $2$-edge cover $H$ is canonical if the following properties are satisfied:
  • For every complex component, each of its pendant blocks contains at least $6$ edges and at least $6$ vertices and each of its non-pendant blocks contains at least $4$ edges and at least $4$ vertices.
Note that a canonical 2-edge cover is also a triangle-free 2-edge cover. We establish the following lemma in Appendix˜B that shows how to convert a minimum triangle-free 2-edge cover of a structured graph into a minimum canonical 2-edge cover in polynomial time.
@@
lemmalemmaCanonicalMain Given a structured graph $G$, in polynomial time we can compute a canonical $2$-edge cover $H$ of $G$, where the number of edges in $H$ is at most (and hence equal to) the number of edges in a minimum triangle-free $2$-edge cover. This definition of canonical and the proof of the above lemma are similar to the ones in [GGA23], except that now components with $7$ edges have to be a $\mathcal{C}_{7}$.
@@ 5 Bridge Covering and Credits
We now have a structured graph $G$ and a minimum canonical 2-edge cover $H$ of $G$ such that $|H|\leq\mathsf{opt}(G)$. We will show how to convert $H$ into a 2-ECSS of $G$ by adding a net total of at most $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$ edges to it. There are two reasons why $H$ is not already a 2-ECSS of $G$: First, $H$ might have multiple components. Second, some of these components might have bridges. We will first modify $H$ so that the resulting components do not have any bridges. Then, we will glue the various resulting 2EC components together to get a desired 2-ECSS, which we defer to the next section. Throughout this transformation, we keep the property that the intermediate solution is a canonical 2-edge cover.
To keep track of how many edges we are adding, we will define a credit invariant that is satisfied throughout the transformation and a cost function. We have a total budget of $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$, which we will distribute all over $H$ as credits of different kinds. The cost of $H$ is defined keeping in mind that the cost of a single edge is 1. Then, $\mathrm{cost}(H)$ denotes the total number of edges in $H$ plus the total credits in $H$. Thus, initially the $\mathrm{cost}(H)$ will be at most $\nicefrac{{5}}{{4}}\cdot\mathsf{opt}$, which will follow from Lemmas 7 and 9. At each step of our transformation, we aim to ensure that the cost never increases, so that the final solution will have cost at most $\nicefrac{{5}}{{4}}\cdot\mathsf{opt}$, and we will be done. We now define our credit invariant and the cost function, which is also used in the next section.
@@
We first show that we can distribute our budget of $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$ over our minimum canonical triangle-free 2-edge cover to satisfy the credit invariant. In other words, we assume the credits are there as required by the credit invariant, and show that $\mathrm{cost}(H)\leq\nicefrac{{5}}{{4}}\cdot\lvert H\rvert$.
{restatable}
lemmalemmaInitialCredit Let $H$ be a canonical $2$-edge cover of a graph $G$. Then, we have $\mathrm{cost}(H)\leq\frac{5}{4}|H|$.
@@
We are now ready to state our main lemma for bridge covering, which essentially is already proved in [GGA23]. Its proof is deferred to Appendix˜C.
@@
We will feed our minimum canonical 2-edge cover $H$ to the algorithm in the above lemma to obtain a bridgeless canonical 2-edge cover $H^{\prime}$, and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)\leq\nicefrac{{5}}{{4}}\cdot\mathsf{opt}(G)$. Each component of $H^{\prime}$ now is 2EC and either has $8$ or more edges, in which case it has a credit of $2$, or it is a cycle with $i$ edges ($\mathcal{C}_{i}$) for $4\leq i\leq 7$ with $\nicefrac{{1}}{{4}}\cdot i$ credits. We will use these credits to glue the 2EC components of $H^{\prime}$ together to obtain a desired 2-ECSS of $G$, as explained in the next section.
@@ 6 Gluing
Using the above lemma, the proof of Theorem˜1 follows easily:
Proof of Theorem˜1.
For a structured graph $G$, we first compute a triangle-free 2-edge cover using Lemma˜4, which can be turned into a canonical 2-edge cover $H$ without increasing the number of edges (Definition˜7). As observed before, we have $|H|\leq\mathsf{opt}$. By Definition˜9, $\mathrm{cost}(H)\leq\frac{5}{4}|H|$ and therefore we have that $\mathrm{cost}(H)\leq\frac{5}{4}\mathsf{opt}$. By Definition˜9, we obtain a bridgeless canonical 2-edge cover $H^{\prime}$ such that $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)\leq\frac{5}{4}\mathsf{opt}$. Finally, we apply Lemma˜5 to turn $H^{\prime}$ into a feasible solution $H^{\prime\prime}$ such that $\mathrm{cost}(H^{\prime\prime})\leq\frac{5}{4}\mathsf{opt}+3$. Now $\mathrm{cr}(H^{\prime\prime})=2$ as it will be a single 2EC component with more than $8$ edges. Thus, $\lvert H^{\prime\prime}\rvert\leq\mathrm{cost}(H^{\prime})+3-2\leq\frac{5}{4}\cdot\mathsf{opt}(G)+1\leq(\frac{5}{4}+\frac{\varepsilon}{4})\cdot\mathsf{opt}(G)$, where the last inequality follows from $\mathsf{opt}(G)\geq\lvert V(G)\rvert\geq\frac{4}{\varepsilon}$, which holds for structured graphs $G$. Now, using the approximation preserving reduction (Definition˜5) for general graphs $G$, we incur an additional loss of $4\varepsilon$. Thus, for any $\varepsilon^{\prime}>0$, we get an approximation ratio of $\frac{5}{4}+\frac{\varepsilon}{4}+4\varepsilon=\frac{5}{4}+\frac{17}{4}\varepsilon=\frac{5}{4}+\varepsilon^{\prime}$ by setting $\varepsilon=\frac{4}{17}\varepsilon^{\prime}$, which proves Theorem˜1. ∎
It remains to prove Lemma˜5. To this end, we first introduce some definitions. In what follows, unless otherwise mentioned, $G$ is a structured graph and $H$ is a bridgeless canonical 2-edge cover of $G$.
@@
The component graph $\bm{\hat{}}{G}_{H}$ w.r.t. $G$ and $H$ is the multigraph obtained from $G$ by contracting the vertices of each 2EC component of $H$ into a single node. We call the vertices of $\bm{\hat{}}{G}_{H}$ nodes to distinguish them from the vertices in $G$. There is a one-to-one correspondence between the edges of $\bm{\hat{}}{G}_{H}$ and $G$. Furthermore, there is a one-to-one correspondence between the nodes of $\bm{\hat{}}{G}_{H}$ and the components of $H$. We will denote the component corresponding to a node $A$ by $C_{A}$.
@@
A non-trivial segment [^2] In the graph theory literature, blocks often refer to maximal 2VC subgraphs, but in this work we use the term blocks to refer to maximal 2EC subgraphs in line with previous works on 2-edge connectivity. Thus, 2VC blocks consisting of at least 3 vertices and non-trivial segments defined above are precisely the same objects. $S$ of $\bm{\hat{}}{G}_{H}$ is a maximal 2-node-connected subgraph of $\bm{\hat{}}{G}_{H}$ consisting of at least 3 nodes. A trivial segment is a single node of $\bm{\hat{}}{G}_{H}$ that is not part of any non-trivial segment. For a segment $S$, $V(S)$ refers to the set of all the vertices in the various components corresponding to the nodes contained in $S$.
Notice that we can break the component graph into segments such that two segments can potentially overlap only at a cut node (follows from Algorithm 4.1.23 in [Wes01]; alternatively see [KV02]).
Since a trivial segment $A$ is not part of any cycles of length at least 3 in $\bm{\hat{}}{G}_{H}$ (otherwise $A$ would be part of a non-trivial segment), $A$ has to be either a cut node of $\bm{\hat{}}{G}_{H}$ or a pendant node (neighbor of exactly one other node) of $\bm{\hat{}}{G}_{H}$. In either case, we can apply the 3-matching lemma (Lemma˜2) to argue that there is a matching of size 3 between $V(C_{A})$ and $V(C_{B})$, where node $B$ is a neighbor of node $A$ in $\bm{\hat{}}{G}_{H}$. The condition of the lemma is satisfied by partitioning $V(G)$ into two parts such that one part has vertices from $V(C_{A})$ and the other part has vertices from $V(C_{B})$ and the only edges going across the parts are between $V(C_{A})$ and $V(C_{B})$. We have the following propositions.
@@
Let $A$ be a trivial segment of $\bm{\hat{}}{G}_{H}$. Then, $A$ is either a cut node of $\bm{\hat{}}{G}_{H}$ or it is a pendant node of $\bm{\hat{}}{G}_{H}$.
@@
Let $A$ and $B$ be adjacent nodes in $\bm{\hat{}}{G}_{H}$ such that $A$ is a trivial segment. Then, there exists a matching of size $3$ going between $V(C_{A})$ and $V(C_{B})$ in $G$.
Having set up the requisite definitions, we now outline the steps involved in the gluing process. First, if $H$ does not contain a huge component, then we first glue a few components together, so that we obtain a canonical bridgeless 2-edge cover $H^{\prime}$ that has a huge component, while maintaining the credit invariant. We will show that such an $H^{\prime}$ can be obtained with $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)+3$. After this step, we assume $H$ has a huge component, an invariant that will be maintained at each iteration of our algorithm. Next, let $L$ be a huge component in $H$. We will distinguish two cases: either $L$ is a trivial segment of $\bm{\hat{}}{G}_{H}$, or it is part of a non-trivial segment of $\bm{\hat{}}{G}_{H}$. In either case, we show that we can make progress, i.e., we can find in polynomial time a canonical bridgeless 2-edge cover $H^{\prime}$ that has fewer components than $H$ such that $H^{\prime}$ contains a huge component and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$. Repeatedly applying the above step will result in a single bridgeless 2EC component, a 2-ECSS of $G$. Thus, if we can establish the existence of these steps, Lemma˜5 follows immediately. These steps are encapsulated in the following three lemmas.
@@
Let $C_{L}$ be a huge component such that $L$ is a trivial segment of $\bm{\hat{}}{G}_{H}$ and $H\neq C_{L}$. Then, in polynomial time we can compute a canonical bridgeless 2-edge cover $H^{\prime}$ of $G$ such that $H^{\prime}$ has fewer components than $H$, $H^{\prime}$ contains a huge component, and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$.
@@
Let $C_{L}$ be a huge component such that $L$ is part of a non-trivial segment of $\bm{\hat{}}{G}_{H}$. Then, in polynomial time we can compute a canonical bridgeless 2-edge cover $H^{\prime}$ of $H$ such that $H^{\prime}$ has fewer components than $H$, $H^{\prime}$ contains a huge component, and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$.
As explained before, the above three lemmas immediately imply Lemma˜5. All that remains is to prove the above three lemmas, which we do in the next three subsections.
6.1 Proof of Lemma˜8